# A Concentration Bound for TD(0) with Function Approximation

### v3

@@ 2 TD(0) - Background and Main Result
There exist finite positive constants $c_{1},c_{2},c_{3}$ and $D$ such that for $0<\delta\leq 1,0<\epsilon\leq 1$, $n_{0}>0$ large enough to satisfy $\alpha+a(n_{0})c_{1}<1$, and $n\geq n_{0}$,
@@
    the inequality
    $\|x_{m}-x^{*}\|\leq e^{-(1-\alpha)b_{n_{0}}(m-1)}\epsilon+\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}},\;\forall n_{0}\leq m\leq n,$
    holds with probability exceeding
    $1-2d\sum_{m=n_{0}+1}^{n}e^{-D\delta^{2}/\beta_{n_{0}}(m)}-P(\|x_{n_{0}}-x^{*}\|>\epsilon)$
@@
    In particular,
    $\|x_{m}-x^{*}\|\leq e^{-(1-\alpha)b_{n_{0}}(m-1)}\epsilon+\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}},\;\forall m\geq n_{0},$
    holds with probability exceeding
    $1-2d\sum_{m\geq n_{0}+1}e^{-D\delta^{2}/\beta_{n_{0}}(m)}-P(\|x_{n_{0}}-x^{*}\|>\epsilon)$
@@ 3 Proof of the Main Result
$z_{n+1}-x^{*}$ $=(1-a(n))(z_{n}-x^{*})+a(n)\left(\sum_{i}\pi(i)F(z_{n},i)-x^{*}\right)$
$=(1-a(n))(z_{n}-x^{*})+a(n)\sum_{i}\pi(i)\left(F(z_{n},i)-F(x^{*},i)\right).$
The second equality follows from the fact that $x^{*}$ is a fixed point for $\sum_{i}\pi(i)F(\cdot,i)$. Then,
$\|z_{n+1}-x^{*}\|$ $\leq(1-a(n))\|z_{n}-x^{*}\|+a(n)\|\sum_{i\in S}\pi_{x_{n}}(i)(F(z_{n},i)-F(x^{*},i))\|$
@@
$\|z_{n}-x^{*}\|\leq\psi(n,n_{0})\|x_{n_{0}}-x^{*}\|\leq e^{-(1-\alpha)b_{n_{0}}(n-1)}\|x_{n_{0}}-x^{*}\|,$ (4)
where $\psi(n,m)\coloneqq\prod_{k=m}^{n-1}(1-(1-\alpha)a(k))$ if $n>m$ and $1$ otherwise. This also implies that for all $n\geq n_{0}$
@@
$+\ a(n)\left(\sum_{i}\pi(i)(F(x_{n},i)-F(z_{n},i))\right)$
$+\ a(n)\left(F(x_{n},Y_{n})-\sum_{i}\pi(i)F(x_{n},i)\right).$
@@
$+\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(\sum_{i\in\mathcal{S}}\pi(i)(F(x_{k},i)-F(z_{k},i))\right)$
$+\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(F(x_{k},Y_{k})-\sum_{i}\pi(i)F(x_{k},i)\right).$ (6)
@@
$F_{1}(Y)=\varphi(Y)r(Y)\in\mathcal{R}^{d}\;\;\;\textrm{and}\;\;\;F_{2}(Y)=\left(\gamma\varphi(Y)\sum_{j}p(j|Y)\varphi(j)^{T}-\varphi(Y)\varphi(Y)^{T}\right)\in\mathcal{R}^{d\times d}.$
We define $U:S\mapsto\mathcal{R}^{d}$ to be a solution of the Poisson equation:
$U(i)=F_{1}(i)-\sum_{j}\pi(j)F_{1}(j)+\sum_{j}p(j|i)U(j),\ i\in S.$ (7)
For $i_{0}\in S$, $\tau\coloneqq\min\{n>0:Y_{n}=i_{0}\}$ and $E_{i}[\cdots]=E[\cdots|Y_{0}=i]$, we know that
$U^{\prime}(i)=E_{i}\left[\sum_{m=0}^{\tau-1}(F_{1}(Y_{m})-\sum_{j\in\mathcal{S}}\pi(j)F_{1}(j))\right],\ i\in S,$ (8)
is one particular solution to the Poisson equation (see, e.g., Lemma 4.2 and Theorem 4.2 of Section VI.4, pp. 85-91, of Borkar (1991)). Thus $\|U^{\prime}(i)\|_{\infty}\leq 2\max_{i}\|F_{1}(i)\|_{\infty}E_{i}[\tau]$. For an irreducible Markov chain with a finite state space, $E_{i}[\tau]$ is finite for all $i$ and hence the solution $U^{\prime}(i)$ is bounded for all $i$. For each $\ell$, the Poisson equation specifies $U^{\ell}(\cdot)$ uniquely only up to an additive constant. Along with the additional constraint that $U(i_{0})=0$ for a prescribed $i_{0}\in S$, the system of equations given by (7) has a unique solution. Henceforth $U$ refers to the unique solution of the Poisson equation with $U(i_{0})=0$. Similarly, let $W:S\mapsto\mathcal{R}^{d\times d}$ be a solution of the Poisson equation:
$W(i)=F_{2}(i)-\sum_{j}\pi(j)F_{2}(j)+\sum_{j}p(j|i)W(j),\ i\in S,$ (9)
with the additional constraint that $W(i_{0})=0$ for a prescribed $i_{0}\in S$ as above.
@@
Here $\widetilde{U}_{k+1}$ and $\widetilde{W}_{k+1}x_{k}$ are martingale difference sequences with respect to $\mathcal{F}_{k}$ where $\widetilde{U}_{k+1}=U(Y_{k+1})-\sum_{j}p(j|Y_{k})U(j)$ and $\widetilde{W}_{k+1}=W(Y_{k+1})-\sum_{j}p(j|Y_{k})W(j)$ for $k\geq n_{0}$ and and the zero vector, resp. the zero matrix, otherwise.
@@
$x_{m+1}-z_{m+1}$ $=\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(\sum_{i\in\mathcal{S}}\pi(i)(F(x_{k},i)-F(z_{k},i))\right)$
$\;\;\;+\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(M_{k+1}x_{k}+\widetilde{W}_{k+1}x_{k}+\widetilde{U}_{k+1}\right)+\mu_{m}(n_{0})$
@@
$\|x_{m+1}-z_{m+1}\|$ $\leq\left\|\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(\sum_{i\in\mathcal{S}}\pi(i)(F(x_{k},i)-F(z_{k},i))\right)\right\|$
@@
$+a(n_{0})(c_{2}+c_{1}\|x_{n_{0}}-x^{*}\|).$
Next we obtain a bound on the probability
$P\left(\|x_{m}-x^{*}\|\leq\exp(-(1-\alpha)b_{n_{0}}(m-1))\epsilon+\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}},\;\forall n_{0}\leq m\leq n\right),$
for some $\epsilon>0$ and $\delta>0$ (recall the assumption that $\alpha+a(n_{0})c_{1}<1$). From (4), recall that $\|z_{n}-x^{*}\|\leq\exp(-(1-\alpha)b_{n_{0}}(n-1))\|x_{n_{0}}-x^{*}\|\;a.s.,$ and hence,
$\|x_{n_{0}}-x^{*}\|\leq\epsilon\implies\|z_{m}-x^{*}\|\leq\exp(-(1-\alpha)b_{n_{0}}(m-1))\|x_{n_{0}}-x^{*}\|.$
@@
$\left\{\|x_{n_{0}}-x^{*}\|\leq\epsilon\right\}\bigcap\left\{x^{\prime}_{n}\leq\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right\}$
$\subseteq\left\{\|x_{m}-x^{*}\|\leq\exp(-(1-\alpha)b_{n_{0}}(m-1))\epsilon+\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}},\;\forall n_{0}\leq m\leq n\right\}.$
@@
$P\left(\|x_{m}-x^{*}\|\leq\exp(-(1-\alpha)b_{n_{0}}(m-1))\|x_{n_{0}}-x^{*}\|+\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}},\;\forall n_{0}\leq m\leq n\right)$
$\geq 1-P\left(x^{\prime}_{n}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\bigcup\|x_{n_{0}}-x^{*}\|>\epsilon\right).$
@@
$\mathbb{B}_{m}=\left\{\xi\mid x_{m}^{\prime}(\xi)>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\;\;\bigcup\;\;\|x_{n_{0}}(\xi)-x^{*}\|>\epsilon\right\}.$
Here the notation $x_{m}^{\prime}(\xi)$ and $x_{n_{0}}(\xi)$ highlights the dependence of $x_{m}^{\prime}$ and $x_{n_{0}}$ on the realizations of $x_{0}$ and $\{Y_{k}\}$. Analogous notation is used for other random variables. For $\xi\notin\mathbb{B}_{n-1}$ let us define $\bar{x}_{k,n-1}(\xi)=x_{k}(\xi)$ and $\bar{z}_{k,n-1}(\xi)=z_{k}(\xi)$ for all $k$. For $\xi\in\mathbb{B}_{n-1}$, we define $\bar{x}_{k,n-1}(\xi)=x^{*}$, and $\bar{z}_{k,n-1}(\xi)=x^{*}$ for all $k$. Also, define $\bar{x}_{m,n-1}^{\prime}(\xi)=\sup_{n_{0}\leq k\leq m}\|\bar{x}_{k,n-1}(\xi)-\bar{z}_{k,n-1}(\xi)\|$. Now, adapting the proof of Proposition 34 from Tao and Vu (2015) for our case, we first note that $\xi\notin\mathbb{B}_{n-1}\implies\bar{x}_{n,n-1}^{\prime}(\xi)=x_{n}^{\prime}(\xi)$, which implies
$P(\bar{x}_{n,n-1}^{\prime}(\xi)\neq x_{n}^{\prime}(\xi))\leq P(\mathbb{B}_{n-1}).$
Henceforth we drop $\xi$ for ease of notation, rendering implicit the dependence of all random variables on $\xi$. Then,
$P\left(x^{\prime}_{n}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right)$
$\leq P\left(x^{\prime}_{n}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\;\bigcup\;\|x_{n_{0}}-x^{*}\|>\epsilon\right)$
$\stackrel{{\scriptstyle(a)}}{{\leq}}P\left(\bar{x}^{\prime}_{n,n-1}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\;\bigcup\;\bar{x}^{\prime}_{n,n-1}\neq x^{\prime}_{n}\;\bigcup\;\|x_{n_{0}}-x^{*}\|>\epsilon\right)$
$\stackrel{{\scriptstyle(b)}}{{\leq}}P\left(\bar{x}^{\prime}_{n,n-1}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right)+P\left(\bar{x}_{n,n-1}^{\prime}\neq x_{n}^{\prime}\;\bigcup\;\|x_{n_{0}}-x^{*}\|>\epsilon\right)$
$\stackrel{{\scriptstyle(c)}}{{\leq}}P\left(\bar{x}^{\prime}_{n,n-1}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right)+P(\mathbb{B}_{n-1})$
$=P\left(\bar{x}^{\prime}_{n,n-1}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right)+P\left(x_{n-1}^{\prime}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\;\;\bigcup\;\;\|x_{n_{0}}-x^{*}\|>\epsilon\right).$
@@
$\left\{\bar{x}^{\prime}_{n,n-1}\leq\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right\}\bigcap\{\bar{x}^{\prime}_{n,n-1}=x^{\prime}_{n}\}\subseteq\left\{x^{\prime}_{n}\leq\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right\},$
@@
$\left\{x^{\prime}_{n}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right\}\subseteq\left\{\bar{x}^{\prime}_{n,n-1}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right\}\bigcup\{\bar{x}^{\prime}_{n,n-1}\neq x^{\prime}_{n}\},$
@@
Now we obtain a bound for $P\left(\bar{x}^{\prime}_{n,n-1}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right)$ by induction. Note that while $\bar{x}^{\prime}_{n-1,n-1}\leq\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}$, $\bar{x}^{\prime}_{n,n-1}$ is not necessarily less than $\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}$. Hence when quantifying the above probability, we only have to look at $\|\bar{x}_{n,n-1}-\bar{z}_{n,n-1}\|$. Then,
$P\left(\bar{x}^{\prime}_{n,n-1}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right)=P\left(\|\bar{x}_{n,n-1}-\bar{z}_{n,n-1}\|>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right).$
We can use (3) to obtain a bound on this probability. In the case where $\xi_{n-1}\in\mathbb{B}_{n-1}$, $\|\bar{x}_{n,n-1}-\bar{z}_{n,n-1}\|=0$ and in the case where $\xi_{n-1}\notin\mathbb{B}_{n-1}$, we have $x_{k}=\bar{x}_{k,n-1}$. So (3) can be written as:

### v4

@@ 3 Main Result
There exist finite positive constants $c_{1},c_{2}$ and $D$ such that for $0<\delta\leq 1,0<\epsilon\leq 1$, $n_{0}>0$ large enough to satisfy $\alpha+a(n_{0})c_{1}<1$, and $n\geq n_{0}$,
@@
    the inequality x_m-x^*∥≤e^-(1-α)b_n_0(m-1)ϵ+a(n0)(c2+c1ϵ)+δ1-α-a(n0)c1,   n_0mn, holds with probability exceeding 1-2d_m=n_0+1^n e^-Dδ^2/β_n_0(m-1)-P(x_n_0-x^*¿ϵ).
@@
    In particular, x_m-x^*∥≤e^-(1-α)b_n_0(m-1)ϵ+a(n0)(c2+c1ϵ)+δ1-α-a(n0)c1,   mn_0, holds with probability exceeding 1-2d_mn_0+1 e^-Dδ^2/β_n_0(m-1n)-P(x_n_0-x^*¿ϵ).
@@ 4 Proof of the Main Result
$z_{n+1}-x^{*}$ $=(1-a(n))(z_{n}-x^{*})+a(n)\left(\sum_{s\in\mathcal{S}}\pi(s)F(z_{n},s)-x^{*}\right)$
$=(1-a(n))(z_{n}-x^{*})+a(n)\sum_{s\in\mathcal{S}}\pi(s)\left(F(z_{n},s)-F(x^{*},s)\right).$
The second equality follows from the fact that $x^{*}$ is a fixed point for $\sum_{s\in\mathcal{S}}\pi(s)F(\cdot,s)$. Then,
$\|z_{n+1}-x^{*}\|$ $\leq(1-a(n))\|z_{n}-x^{*}\|+a(n)\|\sum_{s\in\mathcal{S}}\pi(s)(F(z_{n},s)-F(x^{*},s))\|$
@@
$\|z_{n}-x^{*}\|\leq\prod_{k=n_{0}}^{n-1}(1-(1-\alpha)a(k))\|x_{n_{0}}-x^{*}\|\leq e^{-(1-\alpha)b_{n_{0}}(n-1)}\|x_{n_{0}}-x^{*}\|,$ (4)
This also implies that for all $n\geq n_{0}$,
@@
$+\ a(n)\left(\sum_{s\in\mathcal{S}}\pi(s)(F(x_{n},s)-F(z_{n},s))\right)$
$+\ a(n)\left(F(x_{n},Y_{n})-\sum_{s\in\mathcal{S}}\pi(s)F(x_{n},s)\right).$
@@
$+\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(\sum_{s\in\mathcal{S}}\pi(s)(F(x_{k},s)-F(z_{k},s))\right)$
$+\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(F(x_{k},Y_{k})-\sum_{s\in\mathcal{S}}\pi(s)F(x_{k},s)\right).$ (6)
@@
$F_{1}(Y)=\varphi(Y)r(Y)\in\mathbb{R}^{d}\;\;\;\textrm{and}\;\;\;F_{2}(Y)=\left(\gamma\varphi(Y)\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|Y)\varphi(s^{\prime})^{T}-\varphi(Y)\varphi(Y)^{T}\right)\in\mathbb{R}^{d\times d}.$
We define $U:\mathcal{S}\mapsto\mathbb{R}^{d}$ to be a solution of the Poisson equation:
$U(s)=F_{1}(s)-\sum_{s^{\prime}\in\mathcal{S}}\pi(s^{\prime})F_{1}(s^{\prime})+\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|s)U(s^{\prime}),\ s\in\mathcal{S}.$ (7)
For $s_{0}\in\mathcal{S}$, $\tau\coloneqq\min\{n>0:Y_{n}=s_{0}\}$ and $E_{s}[\cdots]=E[\cdots|Y_{0}=s]$, we know that
$U^{\prime}(s)=E_{s}\left[\sum_{m=0}^{\tau-1}(F_{1}(Y_{m})-\sum_{s^{\prime}\in\mathcal{S}}\pi(s^{\prime})F_{1}(s^{\prime}))\right],\ s\in S,$ (8)
is one particular solution to the Poisson equation (see, e.g., Lemma 4.2 and Theorem 4.2 of Section VI.4, pp. 85-91, of Borkar (1991)). Thus $\|U^{\prime}(s)\|_{\infty}\leq 2\max_{s\in\mathcal{S}}\|F_{1}(s)\|_{\infty}E_{s}[\tau]$. For an irreducible Markov chain with a finite state space, $E_{s}[\tau]$ is finite for all $s$ and hence the solution $U^{\prime}(s)$ is bounded for all $s$. For each $\ell$, the Poisson equation specifies $U^{\ell}(\cdot)$ uniquely only up to an additive constant. Along with the additional constraint that $U(s_{0})=0$ for a prescribed $s_{0}\in S$, the system of equations given by (7) has a unique solution. Henceforth $U$ refers to the unique solution of the Poisson equation with $U(s_{0})=0$. Similarly, let $W:\mathcal{S}\mapsto\mathbb{R}^{d\times d}$ be the unique solution of the Poisson equation:
$W(s)=F_{2}(s)-\sum_{s^{\prime}}\pi(s^{\prime})F_{2}(s^{\prime})+\sum_{s^{\prime}}p(s^{\prime}|s)W(s),\ s\in S,$ (9)
with the additional constraint that $W(s_{0})=0$ for a prescribed $s_{0}\in S$ as above.
@@
Here $\widetilde{U}_{k+1}$ and $\widetilde{W}_{k+1}x_{k}$ are martingale difference sequences with respect to $\mathcal{F}_{k}$ where $\widetilde{U}_{k+1}=U(Y_{k+1})-\sum_{s^{\prime}}p(s^{\prime}|Y_{k})U(s^{\prime})$ and $\widetilde{W}_{k+1}=W(Y_{k+1})-\sum_{s^{\prime}}p(s^{\prime}|Y_{k})W(s^{\prime})$ for $k\geq n_{0}$ and the zero vector, resp. the zero matrix, otherwise.
@@
$x_{m+1}-z_{m+1}$ $=\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(\sum_{s\in\mathcal{S}}\pi(s)(F(x_{k},s)-F(z_{k},s))\right)$
$\;\;\;+\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(M_{k+1}x_{k}+\widetilde{W}_{k+1}x_{k}+\widetilde{U}_{k+1}\right)+\mu_{m}(n_{0}).$
@@
$\|x_{m+1}-z_{m+1}\|$ $\leq\left\|\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(\sum_{s\in\mathcal{S}}\pi(s)(F(x_{k},s)-F(z_{k},s))\right)\right\|$
@@
$+a(n_{0})(c_{2}+c_{1}\|x_{n_{0}}-x^{*}\|).$ (11)
Next we wish to obtain a bound on the probability
$P\Big(\|x_{m}-x^{*}\|\leq\exp(-(1-\alpha)b_{n_{0}}(m-1))\epsilon+\Delta(n_{0},\epsilon,\delta),\;\forall n_{0}\leq m\leq n\Big),$
for some $\epsilon>0$ and $\delta>0$ (recall the assumption that $\alpha+a(n_{0})c_{1}<1$). For ease of notation, here we have defined
$\Delta(n_{0},\epsilon,\delta)\coloneqq\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}.$
From (4), recall that $\|z_{n}-x^{*}\|\leq\exp(-(1-\alpha)b_{n_{0}}(n-1))\|x_{n_{0}}-x^{*}\|\;a.s.,$ and hence,
$\|x_{n_{0}}-x^{*}\|\leq\epsilon\implies\|z_{m}-x^{*}\|\leq\exp(-(1-\alpha)b_{n_{0}}(m-1))\epsilon.$
@@
$\left\{\|x_{n_{0}}-x^{*}\|\leq\epsilon\right\}\bigcap\left\{x^{\prime}_{n}\leq\Delta(n_{0},\epsilon,\delta)\right\}$
$\subseteq\left\{\|x_{m}-x^{*}\|\leq\exp(-(1-\alpha)b_{n_{0}}(m-1))\epsilon+\Delta(n_{0},\epsilon,\delta),\;\forall n_{0}\leq m\leq n\right\}.$
@@
$P\Big(\|x_{m}-x^{*}\|\leq\exp(-(1-\alpha)b_{n_{0}}(m-1))\|x_{n_{0}}-x^{*}\|+\Delta(n_{0},\epsilon,\delta),\;\forall n_{0}\leq m\leq n\Big)$
$\geq 1-P\left(\Big\{x^{\prime}_{n}>\Delta(n_{0},\epsilon,\delta)\Big\}\bigcup\Big\{\|x_{n_{0}}-x^{*}\|>\epsilon\Big\}\right).$
@@
$\mathbb{B}_{m}=\left\{\xi\mid x_{m}^{\prime}(\xi)>\Delta(n_{0},\epsilon,\delta)\;\;\bigcup\;\;\|x_{n_{0}}(\xi)-x^{*}\|>\epsilon\right\}.$
Here the notation $x_{m}^{\prime}(\xi)$ and $x_{n_{0}}(\xi)$ highlights the dependence of $x_{m}^{\prime}$ and $x_{n_{0}}$ on the realizations of $x_{0}$ and $\{Y_{k}\}$. Analogous notation is used for other random variables. For $\xi\notin\mathbb{B}_{n-1}$ let us define $\bar{x}_{k,n-1}(\xi)=x_{k}(\xi)$ and $\bar{z}_{k,n-1}(\xi)=z_{k}(\xi)$ for all $k$. For $\xi\in\mathbb{B}_{n-1}$, we define $\bar{x}_{k,n-1}(\xi)=x^{*}$, and $\bar{z}_{k,n-1}(\xi)=x^{*}$ for all $k$. Also, define $\bar{x}_{m,n-1}^{\prime}(\xi)=\sup_{n_{0}\leq k\leq m}\|\bar{x}_{k,n-1}(\xi)-\bar{z}_{k,n-1}(\xi)\|$. Note that, $\bar{x}^{\prime}_{m,n-1}=0$ when $\xi\in\mathbb{B}_{n-1}$, and $\bar{x}^{\prime}_{m,n-1}=x^{\prime}_{m}\leq\Delta(n_{0},\epsilon,\delta)$ when $\xi\notin\mathbb{B}_{n-1}$. The intuition behind these definitions is that $\bar{x}^{\prime}_{m,n-1}$ is always bounded by $\Delta(n_{0},\epsilon,\delta)$ for all $m\leq n-1$.
Note that $\xi\notin\mathbb{B}_{n-1}\implies\bar{x}_{n,n-1}^{\prime}(\xi)=x_{n}^{\prime}(\xi)$, which implies $P(\bar{x}_{n,n-1}^{\prime}(\xi)\neq x_{n}^{\prime}(\xi))\leq P(\mathbb{B}_{n-1}).$ Henceforth we drop $\xi$ for ease of notation, rendering implicit the dependence of all random variables on $\xi$. Then,
$P\left(x^{\prime}_{n}>\Delta(n_{0},\epsilon,\delta)\right)$ $\leq P\left(x^{\prime}_{n}>\Delta(n_{0},\epsilon,\delta)\;\bigcup\;\|x_{n_{0}}-x^{*}\|>\epsilon\right)$
$\stackrel{{\scriptstyle(a)}}{{\leq}}P\left(\bar{x}^{\prime}_{n,n-1}>\Delta(n_{0},\epsilon,\delta)\;\bigcup\;\bar{x}^{\prime}_{n,n-1}\neq x^{\prime}_{n}\;\bigcup\;\|x_{n_{0}}-x^{*}\|>\epsilon\right)$
$\stackrel{{\scriptstyle(b)}}{{\leq}}P\left(\bar{x}^{\prime}_{n,n-1}>\Delta(n_{0},\epsilon,\delta)\right)+P\left(\bar{x}_{n,n-1}^{\prime}\neq x_{n}^{\prime}\;\bigcup\;\|x_{n_{0}}-x^{*}\|>\epsilon\right)$
$\stackrel{{\scriptstyle(c)}}{{\leq}}P\left(\bar{x}^{\prime}_{n,n-1}>\Delta(n_{0},\epsilon,\delta)\right)+P(\mathbb{B}_{n-1})$
$=P\left(\bar{x}^{\prime}_{n,n-1}>\Delta(n_{0},\epsilon,\delta)\right)+P\left(x_{n-1}^{\prime}>\Delta(n_{0},\epsilon,\delta)\;\;\bigcup\;\;\|x_{n_{0}}-x^{*}\|>\epsilon\right).$
@@
$\left\{\bar{x}^{\prime}_{n,n-1}\leq\Delta(n_{0},\epsilon,\delta)\right\}\bigcap\{\bar{x}^{\prime}_{n,n-1}=x^{\prime}_{n}\}\subseteq\left\{x^{\prime}_{n}\leq\Delta(n_{0},\epsilon,\delta)\right\},$
@@
$\left\{x^{\prime}_{n}>\Delta(n_{0},\epsilon,\delta)\right\}\subseteq\left\{\bar{x}^{\prime}_{n,n-1}>\Delta(n_{0},\epsilon,\delta)\right\}\bigcup\{\bar{x}^{\prime}_{n,n-1}\neq x^{\prime}_{n}\},$
@@
Now we obtain a bound for $P\left(\bar{x}^{\prime}_{n,n-1}>\Delta(n_{0},\epsilon,\delta)\right)$ by induction. We first note that $\bar{x}^{\prime}_{n-1,n-1}$ is bounded by $\Delta(n_{0},\epsilon,\delta)$ by definition. Hence $P\left(\bar{x}^{\prime}_{n,n-1}>\Delta(n_{0},\epsilon,\delta)\right)=P\left(\|\bar{x}_{n,n-1}-\bar{z}_{n,n-1}\|>\Delta(n_{0},\epsilon,\delta)\right)$. We first restate (4) for m=n1m=n-1.
$\|x_{n}-z_{n}\|\leq x^{\prime}_{n}\leq$ $\;(\alpha+a(n_{0})c_{1})x^{\prime}_{n-1}+a(n_{0})(c_{2}+c_{1}\|x_{n_{0}}-x^{*}\|)$
$+\left\|\sum_{k=n_{0}}^{n-1}\chi(m,k+1)a(k)\left(M_{k+1}x_{k}+\widetilde{W}_{k+1}x_{k}+\widetilde{U}_{k+1}\right)\right\|.$
Now, let $I\{\cdot\}$ denote the indicator function which is $1$ when $\{\cdot\}$ holds true, and zero otherwise.
$\|\bar{x}_{n,n-1}-\bar{z}_{n,n-1}\|$
$=\|\bar{x}_{n,n-1}-\bar{z}_{n,n-1}\|I\{\xi_{n-1}\in\mathbb{B}_{n-1}\}+\|\bar{x}_{n,n-1}-\bar{z}_{n,n-1}\|I\{\xi_{n-1}\notin\mathbb{B}_{n-1}\}$
$\stackrel{{\scriptstyle(a)}}{{=}}0\times I\{\xi_{n-1}\in\mathbb{B}_{n-1}\}+\|x_{n}-z_{n}\|\times I\{\xi_{n-1}\notin\mathbb{B}_{n-1}\}$
$\leq I\{\xi_{n-1}\notin\mathbb{B}_{n-1}\}\times\Bigg((\alpha+a(n_{0})c_{1})x^{\prime}_{n-1}+a(n_{0})(c_{2}+c_{1}\|x_{n_{0}}-x^{*}\|)$
$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;+\left\|\sum_{k=n_{0}}^{n-1}\chi(m,k+1)a(k)\left(M_{k+1}x_{k}+\widetilde{W}_{k+1}x_{k}+\widetilde{U}_{k+1}\right)\right\|\Bigg)$
$\stackrel{{\scriptstyle(b)}}{{\leq}}I\{\xi_{n-1}\notin\mathbb{B}_{n-1}\}\times\Bigg((\alpha+a(n_{0})c_{1})\Delta(n_{0},\epsilon,\delta)+a(n_{0})(c_{2}+c_{1}\epsilon)$
$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;+\left\|\sum_{k=n_{0}}^{n-1}\chi(n-1,k+1)a(k)\left(M_{k+1}\bar{x}_{k,n-1}+\widetilde{W}_{k+1}\bar{x}_{k,n-1}+\widetilde{U}_{k+1}\right)\right\|\Bigg).$
Here inequality (a) follows from our definition of $\mathbb{B}_{n-1}$ that $\|\bar{x}_{n,n-1}-\bar{z}_{n,n-1}\|=0$ when $\xi_{n-1}\in\mathbb{B}_{n-1}$, and $x_{n}=\bar{x}_{n,n-1}$ when $\xi_{n-1}\notin\mathbb{B}_{n-1}$. Inequality (b) follows from the fact that when $\xi_{n-1}\notin\mathbb{B}_{n-1}$, then $x_{k}=\bar{x}_{k,n-1}$ for all $k$, and $\|x_{n_{0}}-x^{*}\|\leq\epsilon$. Substituting the expression for $\Delta(n_{0},\epsilon,\delta)$ we obtain the following.
@@
$\;\;\;+\left\|\sum_{k=n_{0}}^{n-1}\chi(n-1,k+1)a(k)\left(M_{k+1}\bar{x}_{k,n-1}+\widetilde{W}_{k+1}\bar{x}_{k,n-1}+\widetilde{U}_{k+1}\right)\right\|.$